# Dust destruction by the supernova remnant forward shock in a turbulent interstellar medium

### v1

@@ 5 Conclusion
We performed 3D MHD Arepo simulations of SNRs evolving in a turbulent ISM that was driven to resemble observations better than previous studies. Different turbulent ISM environments were considered with average ISM densities of $n=1,10,\text{ and }100\,\mathrm{cm^{-3}}$ and turbulent Mach numbers of $\mathcal{M}=0,0.3,1,\text{ and }3$. We let SNRs evolve for the most destructive first 10 kyr and post-processed these simulations with Paperboats to study the dust destruction by the SNR forward shock, considering either purely carbonaceous or purely silicate dust. In detail, we find that:
@@
    The dust destruction by SNR forward shocks in their first 10 kyr of evolution varies significantly, depending on the ISM environment. For n=1100n=1-100 cm-3 and $\mathcal{M}=0-3$, the destruction ranges from 27–92%, which refers to 0.85–11 M of dust. The largest dust mass is destroyed in the high density environment. The least dust is destroyed in the low density ISM where the shock still has end velocities of 250–400 km s-1. Since the total destroyed dust mass has not plateaued yet, it can be expected that more dust will be destroyed in the following years of SNR evolution, as is also derived from comparisons to other studies. Thus, we find that SNRs destroy more ISM dust than they produce and are net dust destroyers.
@@
    Filaments that closely resemble observations slow down the forward shock and therefore shield some dust from destruction, similar to the conclusions of other studies with different inhomogeneities or turbulence driving methods. While filaments in the low density ISM only shield 8–14% of the dust in the first 10 kyr of SNR evolution, the intermediate density cases result in 31–34% less destruction in the highest Mach number simulations compared to the homogeneous ISM. In the high density medium, the forward shock evolves into a thin and dense shell so that grain-grain collisions become significantly more important, resulting in a lower shielding ability of 11–14%.
@@
    Silicates are almost always more efficiently destroyed than carbonaceous grains, which is primarily caused by the higher sputtering yields. In a high density ISM, the destruction difference between these two species decreases, partly caused by more effective grain-grain collisions, which are more efficient for carbonaceous grains than for silicates. In the weakly turbulent and homogeneous high density cases, silicates are even more robust than carbonaceous grains.
@@
    The dust destruction rate can vary even for similar turbulent environments with the same statistical parameters (average ISM density and turbulent Mach number). This comes from the fact that the detailed structure of the turbulence changes, which affects the evolution time and shock velocity at which the first filaments are encountered. We can estimate this difference from the investigated three orthogonal planes, which show a difference in dust destruction rate of up to 9.4% in the first 10 kyr.
@@ Appendix A SNR evolution
Figures A, 9, and A show the temporal SNR evolution of all simulations in this study for the low, intermediate, and high density cases, respectively. Their rows refer to the homogeneous, subsonic, transonic and supersonic simulations and their columns to three different time steps. Links to movies of the SNR evolution are given in the caption of the respective evolution figures.
@@
Figure 9: Same as Fig. A but for the $n=10\,\mathrm{cm^{-3}}$ cases. Here, only the $\mathcal{M}=3$ case has a box length of 40 pc, whereas the other Mach number cases have a box length of 20 pc in each direction. Movies of these simulations can be found on https://youtu.be/boCK842iuds.
@@ Appendix B SN explosion close to filaments
Table 3: Same as Table 1 but for the simulation where the SN explodes close to filaments. [^5] This simulation considers an ISM density of $n=100\,\mathrm{cm^{-3}}$ and a turbulent Mach number of $\mathcal{M}=3$. The dust simulations with grain-grain collisions and sputtering (upper panel) are compared to the simulations that include only sputtering (lower panel). The dust destruction at the end of our simulations (10 kyr, first line) and at similar average shock velocities ($\sim$1 kyr, second line) to all other cases (Appendix F) are shown.
@@
Figure 12: Dust destruction of the $z=0$ slice for the case in which the SN explodes close to filaments (left) and for the case in which the SN explodes in a void (right, also shown in Fig 4). Both simulations have an ISM density of $n=100\,\mathrm{cm^{-3}}$ and Mach number of $\mathcal{M}=3$, and include either silicate (Sil) or carbonaceous (Car) grains, and grain-grain collisions with sputtering (gg+sp) or only sputtering (sp).
@@
The position of the SN in the ISM has a significant impact on its dust destruction. To make all simulations comparable and to mimic pre-SN feedback, we always set off the SN in a large void of the driven turbulence. To quantify the impact of the explosion position, we ran a simulation of the $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ case in which the SN explodes close to filaments, as shown in Fig. 11, and compare it to the fiducial $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ case used in this paper. In this simulation, the SNR only expands up to $\sim$5 pc in radius whereas it expands to $\sim$8 pc in the case where it evolves in a void (bottom panel of Fig. A). The average shock velocity at 10 kyr (see Table 3) is roughly the same between both cases. However, the SNR in a void sweeps up more gas during its evolution because it only encounters dense filaments at the end of the simulation, after it already expanded through a substantial amount of material. The dense filaments can then almost stop the shock since it was continuously slowed down by lower density regions at earlier times. The SNR close to filaments is being slowed down early due to the high gas densities around it. At such early stage, its forward shock is energetic enough to penetrate the filaments and keep slowly expanding.
@@
This difference in SNR evolution affects the dust destruction efficiency. The SNR that explodes close to filaments encounters more dust in its early evolutionary stage where its forward shock is highly energetic. Thus, the destroyed dust mass increases rapidly at the beginning of the simulation, as shown in Fig. 12 but lowers towards the end where the shock is much slower and does not encounter as much dust anymore. The silicate and carbonaceous grains simulations show similar total dust destruction efficiencies of $\sim$86% (Table 3), which is more than 10% more than for the SN explosion in a void (Table 1, 76% and 73% for the silicate and carbonaceous case, respectively). The total destroyed dust mass after 10 kyr is, however, similar because the SNR in a void encounters more material.
@@
When the SNR evolves closer to filaments, it furthermore forms a shell at earlier times since radiative cooling quickly becomes effective due to the high masses of swept-up gas at early times. Therefore, grain-grain collisions become important early on (from $\sim$2–3 kyr onward), resulting in a final difference with the sputtering only simulations of 25% and 33% for silicate and carbonaceous grains, respectively, as shown in Table 3. Grain-grain collisions are also the main reason why carbonaceous grains are similarly well destroyed as silicates in this simulation.
@@
All in all, the dust destruction efficiency of the SNR close to filaments is higher than when it explodes in a void, but similar to the homogeneous and weakly turbulent $n=100\,\mathrm{cm^{-3}}$ cases after 10 kyr of evolution (87% for the silicate and carbonaceous grain simulations). Hence, the filaments cannot shield the dust from destruction at early times if the SN explodes closer to the dense filamentary structures of ISM turbulence but result in one more solar mass of dust that is being destroyed than in the SN in a homogeneous ISM. However, the SNR at filaments is also slowed down quicker than the homogeneous case SNR so that the destruction efficiency and destroyed dust mass are smaller than in the homogeneous case at similar end shock velocities (see Appendix F). After 1 Myr of SNR evolution, the SNR close to filaments might therefore result in lower dust destruction than for the homogeneous high density case.
@@ Appendix C Resolution of the dust simulations
Appendix D Influence of detailed surrounding ISM structure
The SNe in this study are consistently injected in large void regions. The exact distance to nearby filaments and the detailed structure of the ambient ISM, however, still influence the dust destruction by a significant factor because they determine the velocity at which the forward shock encounters high density and dust-rich regions. Our dust simulations with Paperboats only consider three orthogonal 2D slices of the 3D SNR simulations from which we then take the average value for our results. Since the turbulence in these slices has the same statistical properties and the injection region is always a void, the dust destruction difference between these slices can be used to estimate the influence of the detailed structures found in the ambient medium. The difference in dust destruction efficiency between these slices ranges from 0.01% up to 9.4% (Fig. 13), proving that this can have a large impact, especially for the high Mach number simulations since the slice structure can differ more significantly there. The dust destruction difference in one snapshot can increase even more if the SNe were to be injected in regions with higher densities than voids, as shown by Kirchschlager2024a.
@@ Appendix E Grain-size distributions
The grain-size distributions of all simulations are shown in Fig. 14. The black lines show the initial MRN distribution for the carbonaceous and silicate dust, which is altered due to the destruction by the forward shock. The final distributions are similar for the different Mach number cases but differ significantly between the ISM densities. In a low density environment, grain-grain collisions are inefficient. Thus, large grains above $\sim$40 nm mostly survive. However, a large fraction of intermediate sized grains down to 5 nm are sputtered so that the final distribution is altered. Grains that are smaller than 5 nm are not included in the initial distribution but a significant amount is generated by the different destruction mechanisms. The final grain-size distribution of the intermediate density cases are similar but with more destroyed dust. Thus, the difference between initial and final distribution increases. Furthermore, the carbonaceous grain-size distributions show a variation between the different Mach numbers, as large carbonaceous grains are better destroyed in a strongly turbulent medium than in a low Mach number turbulence. Only the high ISM density simulation results significantly deviate from this distribution. There, grain-grain collisions become important because of the formation of the thin and dense shell. Therefore, many large grains are being fragmented to smaller sizes so that intermediate sized grains are less efficiently destroyed than larger grains. Thereby, also much more grains smaller than 5 nm are generated than in the lower density cases. However, if only sputtering is considered, the final grain-sizes are distributed similarly to the intermediate density cases.
@@ Appendix F Dust destruction results with similar forward shock end velocities
Table 4: Same as Table 1 but for a similar average end shock velocity of $387-390\,\mathrm{km\,s^{-1}}$ and thus different end times $t_{\text{end}}$.
@@ Appendix G Total surface area of dust grains
If grain growth of dust grains in the ISM is found to be efficient, the total grain surface area in the post-shock region would be an important parameter to evaluate as these grains could grow in the ISM once they are formed, dependent on their surface area. In all of our parameter sets, we observe that the total number of dust grains significantly increases because grain-grain collisions generate many small grains (see Appendix E). The total grain surface area, however, has a more complex evolution and can decrease if too much dust is being destroyed but also increase in the case that grain-grain collisions are fragmenting enough larger dust grains to smaller sizes with a higher total grain surface area but sputtering is not efficient enough to destroy these grains. Thus, the ratio between the total grain surface area at the end to the initial value at the beginning is dependent on the considered parameter set, as can be seen in Table 5. The total grain surface area of silicate grains was found to always decrease, however, much less for the high density simulations in which grain-grain collisions are efficient. For carbonaceous grains, the total grain surface area increases in the high density cases and the intermediate density case with a strongly turbulent ISM since grain-grain collisions are more efficient for carbonaceous grains than for silicates (see Sect. 3.1). All other parameter sets show a decrease of carbonaceous grain surface area after 10 kyr compared to the initial dust distribution because grain-grain collisions are not that efficient in these simulations. Together with efficient grain growth in the ISM, this could therefore indicate a significantly lower destruction fraction of the SNR forward shock after the most destructive 10 kyr in intermediate and high density cases where grain-grain collisions are important. For some environments, accretion could even be up to almost three times more efficient after the SNR has shocked the ISM than it was before the SN event.
Table 5: Temporal change of the total grain surface area within the SNR shocked region. [^6] The ratio of the total grain surface area at $t=10\,\mathrm{kyr}$ to its initial value at $t=0\,\mathrm{yr}$ is shown, depending on the considered average ISM number density ($n$), turbulent Mach number ($\mathcal{M}$), and grain species (Car for carbonaceous and Sil for silicate grains). A value higher than 1 means that the total grain surface area increased whereas a value lower than 1 means that it decreased.

### v2

@@ 5 Conclusion
We performed 3D MHD Arepo simulations of SNRs evolving in a turbulent ISM that was driven to resemble observations more closely than in previous studies. Different turbulent ISM environments were considered with average ISM densities of $n=1,10,\text{ and }100\,\mathrm{cm^{-3}}$ and turbulent Mach numbers of $\mathcal{M}=0,0.3,1,\text{ and }3$. We let SNRs evolve for the most destructive first 10 kyr and post-processed these simulations with Paperboats to study the dust destruction by the SNR forward shock, considering either purely carbonaceous or purely silicate dust. In detail, we find the following:
@@
    The dust destruction by SNR forward shocks in their first 10 kyr of evolution varies significantly, depending on the ISM environment. For n=1100n=1-100 cm-3 and $\mathcal{M}=0-3$, the destruction ranges from 27–92%, which refers to 0.85–11 M of dust. The largest dust mass is destroyed in the high-density environment. The least dust is destroyed in the low-density ISM where the shock still has end velocities of 250–400 km s-1. Since the total destroyed dust mass has not plateaued yet, it can be expected that more dust will be destroyed in the following years of SNR evolution, as is also derived from comparisons to other studies. Thus, we find that SNRs destroy more ISM dust than they produce, and thus are net dust destroyers.
@@
    Filaments that closely resemble observations slow down the forward shock, and therefore shield some dust from destruction, similarly to the conclusions of other studies with different inhomogeneities or turbulence driving methods. While filaments in the low-density ISM only shield 8–14% of the dust in the first 10 kyr of SNR evolution, the intermediate-density cases result in 31–34% less destruction in the highest Mach number simulations compared to the homogeneous ISM. In the high-density medium, the forward shock evolves into a thin and dense shell so that grain-grain collisions become significantly more important, resulting in a lower shielding ability of 11–14%.
@@
    Silicates are almost always more efficiently destroyed than carbonaceous grains, which is primarily caused by the higher sputtering yields. In a high-density ISM, the destruction difference between these two species decreases, partly caused by more effective grain-grain collisions, which are more efficient for carbonaceous grains than for silicates. In the weakly turbulent and homogeneous high-density cases, silicates are even more robust than carbonaceous grains.
@@
    The dust destruction rate can vary even for similar turbulent environments with the same statistical parameters (average ISM density and turbulent Mach number) because the detailed structure of the turbulence changes, which affects the evolution time and shock velocity at which the first filaments are encountered. We can estimate this difference from the investigated three orthogonal planes, which show a difference in dust destruction rate of up to 9.4% in the first 10 kyr.
@@ Appendix A SNR evolution
Figures A, 9, and A show the temporal SNR evolution of all simulations in this study for the low-, intermediate-, and high-density cases, respectively. Their rows refer to the homogeneous, subsonic, transonic, and supersonic simulations and their columns to three different time steps. Links to movies of the SNR evolution are given in the captions of the respective evolution figures.
@@
Figure 9: Same as Fig. A, but for the $n=10\,\mathrm{cm^{-3}}$ cases. Here, only the $\mathcal{M}=3$ case has a box length of 40 pc, whereas the other Mach number cases have a box length of 20 pc in each direction. Movies of these simulations can be found on https://youtu.be/boCK842iuds.
@@ Appendix B SN explosion close to filaments
Table 3: Same as Table 1, but for the simulation where the SN explodes close to filaments. [^5] This simulation considers an ISM density of $n=100\,\mathrm{cm^{-3}}$ and a turbulent Mach number of $\mathcal{M}=3$. The dust simulations with grain-grain collisions and sputtering (upper panel) are compared to the simulations that include only sputtering (lower panel). The dust destruction at the end of our simulations (10 kyr, first line) and at similar average shock velocities ($\sim$1 kyr, second line) to all other cases (Appendix F) are shown.
@@
Figure 12: Dust destruction of the $z=0$ slice for the case in which the SN explodes close to filaments (left) and for the case in which the SN explodes in a void (right, also shown in Fig. 4). Both simulations have an ISM density of $n=100\,\mathrm{cm^{-3}}$ and Mach number of $\mathcal{M}=3$, and include either silicate (Sil) or carbonaceous (Car) grains, and grain-grain collisions with sputtering (gg+sp) or only sputtering (sp).
@@
The position of the SN in the ISM has a significant impact on its dust destruction. To make all simulations comparable and to mimic pre-SN feedback, we always placed the SN in a large void of the driven turbulence. To quantify the impact of the explosion position, we ran a simulation of the $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ case in which the SN explodes close to filaments, as shown in Fig. 11, and compared it to the fiducial $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ case used in this paper. In this simulation, the SNR only expands up to $\sim$5 pc in radius, whereas it expands to $\sim$8 pc in the case where it evolves in a void (bottom panel of Fig. A). The average shock velocity at 10 kyr (see Table 3) is roughly the same between both cases. However, the SNR in a void sweeps up more gas during its evolution because it only encounters dense filaments at the end of the simulation, after it already expanded through a substantial amount of material. The dense filaments can then almost stop the shock since it was continuously slowed down by lower density regions at earlier times. The SNR close to filaments is slowed down early due to the high gas densities around it. At such early stage, its forward shock is energetic enough to penetrate the filaments and keep slowly expanding.
@@
This difference in SNR evolution affects the dust destruction efficiency. The SNR that explodes close to filaments encounters more dust in its early evolutionary stage where its forward shock is highly energetic. Thus, the destroyed dust mass increases rapidly at the beginning of the simulation, as shown in Fig. 12 but decreases toward the end where the shock is much slower and does not encounter as much dust anymore. The silicate and carbonaceous grains simulations show similar total dust destruction efficiencies of $\sim$86% (Table 3), which is more than 10% more than for the SN explosion in a void (Table 1, 76% and 73% for the silicate and carbonaceous case, respectively). The total destroyed dust mass after 10 kyr is similar, however, because the SNR in a void encounters more material.
@@
When the SNR evolves closer to filaments, it furthermore forms a shell at earlier times since radiative cooling quickly becomes effective due to the high masses of swept-up gas at early times. Therefore, grain-grain collisions become important early on (from $\sim$2–3 kyr onward), resulting in a final difference with the sputtering-only simulations of 25% and 33% for silicate and carbonaceous grains, respectively, as shown in Table 3. Grain-grain collisions are also the main reason why carbonaceous grains are similarly well destroyed as silicates in this simulation.
@@
All in all, the dust destruction efficiency of the SNR close to filaments is higher than when it explodes in a void, but similar to the homogeneous and weakly turbulent $n=100\,\mathrm{cm^{-3}}$ cases after 10 kyr of evolution (87% for the silicate and carbonaceous grain simulations). Hence, the filaments cannot shield the dust from destruction at early times if the SN explodes closer to the dense filamentary structures of ISM turbulence. Instead, they then lead to one more solar mass of destroyed dust than in the SN in a homogeneous ISM. However, the SNR at filaments is also slowed down quicker than the homogeneous case SNR so that the destruction efficiency and destroyed dust mass are smaller than in the homogeneous case at similar end shock velocities (see Appendix F). After 1 Myr of SNR evolution, the SNR close to filaments might therefore result in lower dust destruction than for the homogeneous high-density case.
@@ Appendix C Resolution of the dust simulations
Appendix D Influence of the surrounding ISM structure
The SNe in this study were consistently injected in large void regions. The exact distance to nearby filaments and the detailed structure of the ambient ISM, however, still influence the dust destruction by a significant factor because they determine the velocity at which the forward shock encounters high-density and dust-rich regions. Our dust simulations with Paperboats only consider three orthogonal 2D slices of the 3D SNR simulations from which we then take the average value for our results. Since the turbulence in these slices has the same statistical properties and the injection region is always a void, the dust destruction difference between these slices can be used to estimate the influence of the detailed structures found in the ambient medium. The difference in dust destruction efficiency between these slices ranges from 0.01% up to 9.4% (Fig. 13), proving that this can have a large impact, especially for the high Mach number simulations since the slice structure can differ more significantly there. The dust destruction difference in one snapshot can increase even more if the SNe are injected in regions with higher densities than voids, as shown by Kirchschlager et al. (2024a).
@@ Appendix E Grain-size distributions
The grain-size distributions of all simulations are shown in Fig. 14. The black lines show the initial MRN distribution for the carbonaceous and silicate dust, which is altered due to the destruction by the forward shock. The final distributions are similar for the different Mach number cases but differ significantly between the ISM densities. In a low-density environment, grain-grain collisions are inefficient. Thus, large grains above $\sim$40 nm mostly survive. However, a large fraction of intermediate-sized grains down to 5 nm is sputtered so that the final distribution is altered. Grains that are smaller than 5 nm are not included in the initial distribution but a significant number is generated by the different destruction mechanisms. The final grain-size distributions of the intermediate-density cases are similar but with more destroyed dust. Thus, the difference between initial and final distribution increases. Furthermore, the carbonaceous grain-size distributions show a variation between the different Mach numbers, as large carbonaceous grains are better destroyed in a strongly turbulent medium than in low Mach number turbulence. Only the high-density ISM simulations deviate significantly from this distribution. There, grain-grain collisions become important because of the formation of the thin and dense shell. Therefore, many large grains are fragmented to smaller sizes so that intermediate-sized grains are less efficiently destroyed than larger grains. In addition, many more grains smaller than 5 nm are generated than in the lower density cases. However, if only sputtering is considered, the final grain-sizes are distributed similarly to the intermediate-density cases.
@@ Appendix F Dust destruction results with similar forward shock end velocities
Table 4: Same as Table 1, but for a similar average end shock velocity of $387-390\,\mathrm{km\,s^{-1}}$, and thus different end times $t_{\text{end}}$.
@@ Appendix G Total surface area of dust grains
If grain growth of dust grains in the ISM is found to be efficient, the total grain surface area in the post-shock region is an important parameter to evaluate as these grains could grow in the ISM once they are formed, dependent on their surface area. In all of our parameter sets, we observe that the total number of dust grains significantly increases because grain-grain collisions generate many small grains (see Appendix E). The total grain surface area, however, has a more complex evolution and can decrease if too much dust is destroyed but also increase in the case that grain-grain collisions are fragmenting enough larger dust grains to smaller sizes with a higher total grain surface area but sputtering is not efficient enough to destroy these grains. Thus, the ratio between the total grain surface area at the end to the initial value at the beginning is dependent on the considered parameter set, as can be seen in Table 5. The total grain surface area of silicate grains is always found to decrease; however, it decreases much less for the high-density simulations in which grain-grain collisions are efficient. For carbonaceous grains, the total grain surface area increases in the high-density cases and the intermediate-density case with a strongly turbulent ISM since grain-grain collisions are more efficient for carbonaceous grains than for silicates (see Sect. 3.1). All other parameter sets show a decrease of carbonaceous grain surface area after 10 kyr compared to the initial dust distribution because grain-grain collisions are not that efficient in these simulations. Together with efficient grain growth in the ISM, this could therefore indicate a significantly lower destruction fraction of the SNR forward shock after the most destructive 10 kyr in intermediate- and high-density cases where grain-grain collisions are important. For some environments, accretion could even be up to almost three times more efficient after the SNR shocks the ISM than before the SN event.
Table 5: Temporal change of the total grain surface area within the SNR shocked region. [^6] The ratio of the total grain surface area at $t=10\,\mathrm{kyr}$ to its initial value at $t=0\,\mathrm{yr}$ is shown, depending on the considered average ISM number density ($n$), turbulent Mach number ($\mathcal{M}$), and grain species (Car for carbonaceous and Sil for silicate grains). A value higher than 1 means that the total grain surface area increased, whereas a value lower than 1 means that it decreased.